OpenStudy (anonymous):

find the complex conjugate of (3+4i)^2 ...

OpenStudy (lilg132):

-7 + 24i

OpenStudy (anonymous):

plz explain

OpenStudy (slaaibak):

(3 + 4i)^2 = 9 + 24i + 16i^2 = 9 + 24i - 16 = -7 + 24i therefore the conjugate = -7 - 24i

OpenStudy (anonymous):

24i+7

OpenStudy (slaaibak):

nope, it's: -7 - 24i

OpenStudy (anonymous):

no!!!! its 7+24i

OpenStudy (lilg132):

how did you get 9 + 24i?

OpenStudy (slaaibak):

No. http://en.wikipedia.org/wiki/Complex_conjugate

OpenStudy (anonymous):

k

OpenStudy (anonymous):

hey doesnt the conjugate of -7+24i = 7-24i ?????

OpenStudy (slaaibak):

Nope. In mathematics, complex conjugates are a pair of complex numbers, both having the same real part, but with imaginary parts of equal magnitude and opposite signs.[ -Wikipedia

OpenStudy (wasiqss):

dhash =dumb

OpenStudy (wasiqss):

:P

OpenStudy (anonymous):

ok thanks all